Weather sunny and warm 24 degrees.   Inspections (except #7) not undertaken last week due to poor weather.
Hive #1 – inspected by Rowan.  14×12.  Queen seen Brood on 6 frames and plenty of stores.  Moved empty frame inside pollen barrier.
Hive #2 – inspected by Anne.  Double brood.  Queen seen.  Brood on 7 frames.  Frame of eggs and brood moved to #5.
Hive #3 – inspected by Peter R.  Single Deep.  Queen seen.  Brood on 6 frames.   Brought super below crown board.
Hive #4 – inspected by Hilary.  Double brood.  Queen seen.  Brood on 5 frames in Deep 1 and 6 frames in Deep 2.  Plenty of stores and room.  Chalk brood and sac brood seen – marked affected frame with ‘?’.  Frame 5 in Deep 2 set aside as drone trap.
Hive #5 – inspected by Hilary.  Brood + half.  Queen not seen.  2 charged queen cells (frame marked X) but no brood.   Removed the two queen cells.   Although two frames had been added from #7 two weeks ago, the cells are likely to be from the failed queen rather than the new frames.   Further frame added this week from #2.   If this fails then consider combining the colony.
Hive #6 – inspected by ?   Double brood. Queen not seen.  9 frames of brood in top Deep, 7 in bottom Deep. Plenty of stores and room.  High proportion of drone brood in bottom Deep.  Consider removing next week.
Hive #7 – inspected by Brian. Double brood.  Queen seen.  14 frames of brood – 10 in bottom Deep and 4 in top Deep.  Plenty of room and stores.  Foundation added last week had been drawn on 4 frames.   1:1 syrup added.
Hive #8 – inspected by Liz.  Double brood.  Queen not seen.  5 frames of brood on Deep 2, 2 frames on Deep 1.   Drone brood on Deep 2 removed.  Consider splitting next week.